# MetaSym: A Symplectic Meta-learning Framework for Physical Intelligence

### v1

@@ 4.4 Benchmarks
We evaluate our architecture’s accuracy and memory footprint against two state-of-the-art methods and a naive baseline: an auto-regressive Transformer model (Vaswani, 2017), a physics-informed DHNN model (Greydanus et al., 2019) and a Multilayered Perceptron (MLP). Both MetaSym and the Transformer predict the next timestep autoregressively, while the DHNN estimates the symplectic and dissipative gradients and relies on an integrator to unroll the trajectory. Consequently, the DHNN is also unable to predict quadrature measurements since these derivatives are not well defined and involve a complex Wiener term (Milburn, 1987). Additionally, the DHNN’s training proved to be less stable as compared to MetaSym and the Transformer’s convergence properties. Finally, all models were tested in an autoregressive manner, in which each prediction is provided again as input to the model for the next timestep. As seen in Table 1, this study exposes the true performance capability and scalability of our architecture with the superior error accumulation over long horizon prediction of MetaSym. We achieve superior performance with smaller model sizes for MetaSym especially for high dimensional systems eg. spring mesh system. For the quantum dynamics benchmark, smaller models are crucial to achieve real-time prediction and control. Hardware limitations for quantum control restrict the use of larger models with higher latency. We demonstrate superior performance compared to the Transformer, the current state-of-the-art model (Vaidhyanathan et al., 2024). Moreover, our model size falls below the threshold for real-time predictions of quantum devices (Reuer et al., 2023).
Table 1: Trajectory error and parameter count against SoTA baselines for three out-of-distribution datasets
@@ 5 Conclusion & Future Work
From a geometric perspective, the SymplecticEncoder approximates canonical flows while preserving key invariants, significantly mitigating energy drift and ensuring robust long-term predictions. The encoder’s invertible design enforces time-reversal symmetry and reduces error accumulation. Meanwhile, the ActiveDecoder models departures from ideal Hamiltonian evolution through autoregressive prediction and meta-attention. The resulting architecture is computationally efficient, given that it does not require explicit numerical integration during inference, and—through meta-learning—readily adapts to system variations with minimal additional data. This approach offers a scalable and unified framework for high-fidelity physics modeling in realistic settings with provable near symplectic properties (see Appendix C.4).
Considering MetaSym’s promising performance, we seek to investigate several future directions such as incorporating a fully symplectic network for modeling realistic physics by exploiting the underlying structure of noise and our control signals. Another natural extension of few-shot adaptation is online learning and control. Since MetaSym can quickly adapt to new system configurations with minimal data, this capability can be leveraged in real-time control loops and model-based Reinforcement Learning (RL) algorithms. Future research can investigate how to integrate the decoder’s adaptation step with RL or adaptive Model Predictive Control (MPC) frameworks, effectively enabling self-tuning controllers in rapidly changing environments.
Acknowledgments
The authors would like to thank Lucas Schorling, Federico Fedele and Vivek Wadhia for the useful discussions. P.V. is supported by the United States Army Research Office under Award No. W911NF-21-S-0009-2. A.P is supported by University of Oxford’s Clarendon Fund. M.T.M. is supported by a Royal Society University Research Fellowship. N.A. acknowledges support from the European Research Council (grant agreement 948932) and the Royal Society (URF-R1-191150).
@@ Appendix A Meta Learning Setup
As mentioned in Section 3, we elaborate on the algorithmic pipeline of our SymplecticEncoder and ActiveDecoder modules. In particular, we outline the MAML-based meta-learning strategy for the SymplecticEncoder and the inner-outer adaptation meta-learning for the ActiveDecoder. Figure 5 provides an interpretable insight into the effect of inner adaptation during the meta-update step and its effectiveness.
@@ A.1 Encoder
7:            $\mathcal{I}_{\mathrm{adapt}}\subset\{1,\dots,\mathcal{T}_{i}\},\quad\mathcal{I}_{\mathrm{meta}}=\{1,\dots,\mathcal{T}_{i}\}\setminus\mathcal{I}_{\mathrm{adapt}}$.
@@
15:                 $\bigl{(}\hat{\mathbf{q}}^{(i)}_{t+1},\hat{\mathbf{p}}^{(i)}_{t+1}\bigr{)}\;\leftarrow\;\Phi_{\mathrm{SE}}\!\Bigl{(}\mathbf{q}^{(i)}_{t},\,\mathbf{p}^{(i)}_{t},\,\mathbf{u}^{(i)}_{t};\,\theta_{SE}^{(i)}\Bigr{)}$
@@
18:              $\mathcal{L}_{\mathrm{inner}}^{(i)}.\texttt{backward()}$()
@@
24:              $\bigl{(}\hat{\mathbf{q}}^{(i)}_{t+1},\hat{\mathbf{p}}^{(i)}_{t+1}\bigr{)}\;\leftarrow\;\Phi_{\mathrm{SE}}\!\Bigl{(}\mathbf{q}^{(i)}_{t},\,\mathbf{p}^{(i)}_{t},\,\mathbf{u}^{(i)}_{t};\,\theta_{SE}\Bigr{)}$
@@
28:        $\mathcal{L}_{\mathrm{meta}}.\texttt{backward()}$() {Eq. (5)}
@@ A.2 Decoder
6:            $\mathcal{I}_{\mathrm{adapt}}\subset\{1,\dots,\mathcal{T}_{i}\},\quad\mathcal{I}_{\mathrm{meta}}=\{1,\dots,\mathcal{T}_{i}\}\setminus\mathcal{I}_{\mathrm{adapt}}$.
@@
12:                 $\bigl{(}\hat{\mathbf{q}}^{(i)}_{t+1},\hat{\mathbf{p}}^{(i)}_{t+1}\bigr{)}\;\leftarrow\;\Phi_{\mathrm{AD}}\!\Bigl{(}\mathbf{q}^{(i)}_{t},\,\mathbf{p}^{(i)}_{t},\,\mathbf{u}^{(i)}_{t};\,\theta_{AD},\,\zeta_{i}\Bigr{)}$
@@
14:                 $\mathcal{L}_{\mathrm{inner}}^{(i)}.\texttt{backward()}$() {Eq. (6)}
@@
22:              $\bigl{(}\hat{\mathbf{q}}^{(i)}_{t+1},\hat{\mathbf{p}}^{(i)}_{t+1}\bigr{)}\;\leftarrow\;\Phi_{\mathrm{AD}}\!\Bigl{(}\mathbf{q}^{(i)}_{t},\,\mathbf{p}^{(i)}_{t},\,\mathbf{u}^{(i)}_{t};\,\theta_{AD},\,\zeta_{i}^{*}\Bigr{)}$
@@
25:           $\mathcal{L}_{\mathrm{outer}}^{(i)}.\texttt{backward()}$() {Eq. (7)}
@@
Figure 5: Represents the change in loss between training steps due to the meta-update of parameters.
@@ B.1 Spring Mesh System
Spring networks are a simple proxy for numerous physical scenarios involving deformable solids and cloth (in computer graphics, mechanics, or robotics). Each pair of connected particles exchanges spring forces that depend on displacements from rest lengths. Viscous damping terms further shape the evolution. While the individual force law (Hooke’s law) is straightforward, the combination of hundreds of coupled springs can give rise to complex large-scale deformations and oscillations. By benchmarking on a spring mesh, we can examine how MetaSym learns large deformations, wave propagation through a membrane, or the impact of damping.
The training dataset consists of 25 distinct spring-mesh systems, each characterized by a unique set of physical parameters, including spring stiffness, damping coefficients, and initial conditions as indicated by Table 3. These parameters are sampled from a predefined distribution to ensure sufficient diversity within the training set. Each system is simulated over a time span of 2000 irregular timesteps, capturing the full trajectory of node displacements and momenta. The resulting dataset provides a rich representation of dynamical behaviors within the parameter space.
To assess generalization and robustness, we construct a validation dataset comprising 10 additional spring-mesh systems. Unlike the training set, the parameters for these systems are drawn from distributions that differ from those used during training as indicated by Table 3, introducing a domain shift that mimics real-world variations. This OOD validation set enables a rigorous evaluation of the model’s ability to extrapolate beyond the observed training dynamics and adapt to unseen conditions. For further information regarding the spring-mesh benchmark refer to (Otness et al., 2021).
@@
$\gamma_{decay}\;\left[kg/s\right]$ $\mathrm{Uniform}(0.1,0.2)$
mass $\left[kg\right]$ $\mathrm{Uniform}(0.1,2.0)$
@@
initial perturbations $[m]$ $\mathrm{Uniform}(0,0.6)$
@@
Table 2: In-distribution parameters $T=2000$.
@@
initial perturbations $[m]$ $\mathrm{Uniform}(0.9,2.5)$
@@
Table 3: Out-of-distribution parameters, $T=2000$
@@ B.2 Open Quantum System Derivation
Understanding the behavior of quantum systems plays a vital role in the development of promising technologies such as quantum computing, metrology, and sensing. While deep learning has found great success in several areas such as quantum control (Vaidhyanathan et al., 2024), error correction (Bausch et al., 2024) and tuning (Ares, 2021; Gebhart et al., 2023), predicting the measurement record based on modeling quantum dynamics has long remained elusive. In many scenarios, the system of interest is open: it couples to an environment (or bath) that can introduce thermal noise and dissipation. Furthermore, continuous monitoring (e.g., via homodyne detection) adds additional measurement backaction, reflecting fundamental constraints from quantum mechanics (Jacobs & Steck, 2006). Capturing these noise and measurement effects is pivotal for accurately predicting quantum trajectories and devising robust control protocols.
Unlike closed Hamiltonian evolutions, open quantum systems require one to solve Stochastic Master Equations (SMEs) incorporating decoherence and measurement terms. These equations produce trajectories of the (mixed) quantum state conditioned on the noisy measurement record. In many practical settings, however, we only have direct access to certain observables (e.g., position and momentum quadratures) rather than the full quantum state. Hence, training a deep learning network to model the quantum system and to predict future measurement outcomes becomes a natural and practically relevant challenge. The SME describes the evolution of the conditioned quantum state $\rho_{c}(t)$ under the effect of environmental and measurement noise as (Jacobs & Steck, 2006)
@@
where the dissipator is $\mathcal{D}[\hat{L}]\rho=\hat{L}\rho\hat{L}^{\dagger}-\frac{1}{2}\left\{\hat{L}^{\dagger}\hat{L},\rho\right\}$. Each Linblad operator, $\hat{L}_{j}$, represents a collapse operator that encodes coupling to the environment (e.g., photon loss to a reservoir, thermal excitations, dephasing, etc.) (Manzano, 2020). The stochastic backaction term, $\mathcal{H}[\hat{M}]\rho=\hat{M}\rho+\rho\hat{M}^{\dagger}-\operatorname{Tr}\left[\left(\hat{M}+\hat{M}^{\dagger}\right)\rho\right]\rho$, describes continuous monitoring of an observable $\hat{M}$, with $\eta$ the measurement efficiency ($0\leq\eta\leq 1$). The Wiener increment, $dW_{t}$, captures the randomness inherent in quantum measurement outcomes.
@@ B.2.1 Setup for Parameteric Oscillator
$\hat{L}_{1}=\sqrt{\gamma\left(\bar{n}_{{th}}+1\right)}\hat{\boldsymbol{a}},\quad\hat{L}_{2}=\sqrt{\gamma\bar{n}_{\mathrm{th}}}\hat{a}^{\dagger},$ (10)
@@
We employ heterodyne detection of the field operator $\hat{a}$. Based on this measurement scheme, we can get seperate the real and imaginary parts to obtain quadrature values that roughly correspond to $X$ and $P$ while adding quantum noise and measurement uncertainities due to quantum mechanical effect (Flurin et al., 2020).
In the following Tables 5 and 5, we describe the parameters used to generate our dataset by solving the SME in order to generate training data.
Parameters Values
@@
thermal occupation $\langle n_{th}\rangle$ $\mathrm{Uniform}(0.1,0.5)$
@@
Table 4: In-distribution parameters $dt=0.5$, $T=600$.
Parameters Values
@@
thermal occupation $\langle n_{th}\rangle$ $\mathrm{Uniform}(0.6,0.7)$
@@
Table 5: Out-of-distribution parameters $dt=0.5$, $T=600$

### v2

@@ 4.4 Benchmarks
We benchmarked MetaSym against three baselines: an autoregressive Transformer [18], a physics-informed dissipative Hamiltonian neural network (DHNN) [11], and a naive Multi-Layer Perceptron (MLP). Both MetaSym and the Transformer perform autoregressive next-timestep prediction, whereas the DHNN models the system’s symplectic and dissipative gradients, relying on an integrator for trajectory unrolling. The MLP also predicts the next state directly. Unlike the others, the DHNN cannot predict quadrature measurements due to ill-defined derivatives involving complex Wiener terms [43], and its training proved less stable than that of MetaSym and the Transformer. All models were evaluated in an autoregressive rollout regime. As shown in Table 1, MetaSym consistently outperforms the alternatives in long-horizon prediction accuracy, particularly when using a 30-timestep context window. It achieves lower error accumulation and requires fewer parameters, which is especially advantageous for high-dimensional systems like spring-meshes. The variance present across all models is attributed to the noise that we artificially added to each network’s input in order to highlight MetaSym’s robustness and real-world applicability. In quantum dynamics, where real-time prediction and control are constrained by hardware latency and memory [44], MetaSym’s compact architecture offers a significant advantage over Transformer-based architectures [45].
Table 1: Trajectory error and parameter count against SoTA baselines for three OOD datasets
@@ 5 Conclusion
From a geometric perspective, the SymplecticEncoder approximates canonical flows while preserving key invariants, significantly mitigating energy drift and ensuring robust long-term predictions. The encoder’s invertible design enforces time-reversal symmetry and reduces error accumulation. Meanwhile, the ActiveDecoder models departs from ideal Hamiltonian evolution through autoregressive prediction and meta-attention. The resulting architecture is computationally efficient, given that it does not require explicit numerical integration during inference, and, through meta-learning, it readily adapts to system variations with minimal additional data. This approach offers a scalable and unified framework for high-fidelity physics modeling in realistic settings with provable near-symplectic properties (see Appendix C.4).
Limitations & Future Work: Considering MetaSym’s promising performance, we seek to investigate several future directions such as incorporating a fully symplectic network for modeling realistic physics by exploiting the underlying structure of dissipation and our control signals. MetaSym’s effectiveness on real world system data also remains to be investigated. Another natural extension of few-shot adaptation is online learning and control, since MetaSym can quickly adapt to new system configurations with minimal data. This could be leveraged in real-time control loops and model-based Reinforcement Learning (RL) algorithms. Finally, future research could focus on how to integrate the decoder’s adaptation step with RL or adaptive Model Predictive Control (MPC) frameworks, effectively enabling self-tuning controllers in rapidly changing environments.
Acknowledgements
The authors would like to thank Lucas Schorling, Vassil Atanassov, Federico Fedele and Vivek Wadhia for the useful discussions. P.V. is supported by the United States Army Research Office under Award No. W911NF-21-S-0009-2. A.P is supported by University of Oxford’s Clarendon Fund. M.T.M. is sup- ported by a Royal Society University Research Fellowship. N.A. acknowledges support from the European Research Council (grant agreement 948932) and the Royal Society (URF-R1-191150).
@@ Appendix A Meta Learning Setup
As mentioned in Section 3, we elaborate on the algorithmic pipeline of our SymplecticEncoder and ActiveDecoder modules. In particular, we outline the MAML-based meta-learning strategy for the SymplecticEncoder and the bi-level adaptation meta-learning for the ActiveDecoder. Figure 5 provides an interpretable insight into the effect of inner adaptation during the meta-update step and its effectiveness.
@@ A.1 Encoder
7:            $\mathcal{I}_{\mathrm{adapt}}\subset\{1,\dots,\mathcal{T}^{(i)}_{\mathrm{adapt}}\},\quad\mathcal{I}_{\mathrm{meta}}=\{1,\dots,\mathcal{T}^{(i)}_{\mathrm{meta}}\}\setminus\mathcal{I}_{\mathrm{adapt}}$.
@@
15:                 $\bigl{(}\hat{\mathbf{q}}^{(i)}_{t+1},\hat{\mathbf{p}}^{(i)}_{t+1}\bigr{)}\;\leftarrow\;\Phi_{\theta_{\mathrm{SE}}}\!\Bigl{(}\mathbf{q}^{(i)}_{t},\,\mathbf{p}^{(i)}_{t},\,\mathbf{u}^{(i)}_{t};\,\theta_{SE}^{(i)}\Bigr{)}$
@@
18:              ($\mathcal{L}_{\mathrm{inner}}^{(i)}/\mathcal{T}_{adapt}^{(i)}).\texttt{backward()}$()
@@
24:              $\bigl{(}\hat{\mathbf{q}}^{(i)}_{t+1},\hat{\mathbf{p}}^{(i)}_{t+1}\bigr{)}\;\leftarrow\;\Phi_{\theta_{\mathrm{SE}}}\!\Bigl{(}\mathbf{q}^{(i)}_{t},\,\mathbf{p}^{(i)}_{t},\,\mathbf{u}^{(i)}_{t};\,\theta_{SE}\Bigr{)}$
@@
28:        ($\mathcal{L}_{\mathrm{meta}}/\mathcal{T}_{meta}^{(i)}).\texttt{backward()}$() {Eq. (3)}
@@ A.2 Decoder
6:            $\mathcal{I}_{\mathrm{adapt}}\subset\{1,\dots,\mathcal{T}_{\mathrm{adapt}}^{(i)}\},\quad\mathcal{I}_{\mathrm{meta}}=\{1,\dots,\mathcal{T}_{\mathrm{meta}}^{(i)}\}\setminus\mathcal{I}_{\mathrm{adapt}}$.
@@
12:                 $\bigl{(}\hat{\mathbf{q}}^{(i)}_{t+1},\hat{\mathbf{p}}^{(i)}_{t+1}\bigr{)}\;\leftarrow\;\Phi_{\theta_{\mathrm{AD}}}\!\Bigl{(}\mathbf{q}^{(i)}_{t:c},\,\mathbf{p}^{(i)}_{t:c},\,\mathbf{u}^{(i)}_{t:c};\,\theta_{AD},\,\zeta_{i}\Bigr{)}$
@@
14:                 ($\mathcal{L}_{\mathrm{inner}}^{(i)}/\mathcal{T}_{adapt}^{(i)}).\texttt{backward()}$() {Eq. (4)}
@@
22:              $\bigl{(}\hat{\mathbf{q}}^{(i)}_{t+1},\hat{\mathbf{p}}^{(i)}_{t+1}\bigr{)}\;\leftarrow\;\Phi_{\theta_{\mathrm{AD}}}\!\Bigl{(}\mathbf{q}^{(i)}_{t:c},\,\mathbf{p}^{(i)}_{t:c},\,\mathbf{u}^{(i)}_{t:c};\,\theta_{AD},\,\zeta_{i}^{*}\Bigr{)}$
@@
25:           ($\mathcal{L}_{\mathrm{outer}}^{(i)}/\mathcal{T}_{meta}^{(i)}).\texttt{backward()}$() {Eq. (5)}
@@ B.1 Spring Mesh System
Spring networks are a simple proxy for numerous physical scenarios involving deformable solids and cloth (in computer graphics, mechanics, or robotics). Each pair of connected particles exchanges spring forces that depend on displacements from rest lengths. Viscous damping terms further shape the evolution. While the individual dynamics (Hooke’s law) are straightforward, the combination of hundreds of coupled springs can give rise to complex large-scale deformations and oscillations. By benchmarking on a spring-mesh, we can examine how MetaSym learns large deformations, wave propagation through a membrane, or the impact of damping.
The training dataset consists of 25 distinct spring-mesh systems, each characterized by a unique set of physical parameters, including spring stiffness, damping coefficients, and initial conditions as indicated by Table 2(a). These parameters are sampled from a predefined distribution to ensure sufficient diversity within the training set. Each system is simulated over a time span of 2000 irregular timesteps, capturing the full trajectory of node displacements and momenta. The resulting dataset provides a rich representation of dynamical behaviors within the parameter space.
To assess generalization and robustness, we construct a test dataset comprising ten additional spring-mesh systems. Unlike the training set, the parameters for these systems are drawn from distributions that differ from those used during training as indicated by Table 2(b), introducing a domain shift that mimics real-world variations. This OOD test set enables a rigorous evaluation of the model’s ability to extrapolate beyond the observed training dynamics and adapt to unseen conditions. For further information regarding the spring-mesh benchmark refer to [1].
@@
$\gamma_{decay}$ [kg/s] $\mathrm{Uniform}(0.1,0.2)$
mass [kg] $\mathrm{Uniform}(0.1,2.0)$
@@
Init. Conds. [m] $\mathrm{Uniform}(0,0.6)$
@@
Mesh Size [$n_{x}\times n_{y}$] 10 $\times$ 10
(a) In-distribution parameters $T=2000$
@@
Init. Conds. [m] $\mathrm{Uniform}(0.9,2.5)$
@@
Mesh Size [$n_{x}\times n_{y}$] 10 $\times$ 10
(b) Out-of-distribution parameters $T=2000$
Table 2: Parameter ranges for in-distribution and out-of-distribution regimes.
@@ B.2 Open Quantum System Derivation
Understanding the behavior of quantum systems plays a vital role in the development of promising technologies such as quantum computing, metrology, and sensing. While deep learning has found great success in several areas such as quantum control [45], error correction [47] and tuning [48, 49], predicting the measurement record based on modeling quantum dynamics has long remained elusive. In many scenarios, the system of interest is open: it couples to an environment (or bath) that can introduce thermal noise and dissipation. Furthermore, continuous monitoring (e.g., via homodyne detection) adds additional measurement backaction, reflecting fundamental constraints from quantum mechanics [50]. Capturing these noise and measurement effects is pivotal for accurately predicting quantum trajectories and devising robust control protocols.
Unlike closed Hamiltonian evolutions, open quantum systems require one to solve Stochastic Master Equations (SMEs) incorporating decoherence and measurement terms. These equations produce trajectories of the (mixed) quantum state conditioned on the noisy measurement record. In many practical settings, however, we only have direct access to certain observables (e.g., position and momentum quadratures) rather than the full quantum state. Hence, training a deep learning network to model the quantum system and to predict future measurement outcomes becomes a natural and practically relevant challenge. The SME describes the evolution of the conditioned quantum state $\rho_{c}(t)$ under the effect of environmental and measurement noise as [50]
@@
where the dissipator is $\mathcal{D}[\hat{L}]\rho=\hat{L}\rho\hat{L}^{\dagger}-\frac{1}{2}\left\{\hat{L}^{\dagger}\hat{L},\rho\right\}$. Each Linblad operator, $\hat{L}_{j}$, represents a collapse operator that encodes coupling to the environment (e.g., photon loss to a reservoir, thermal excitations, dephasing, etc.) [51]. The stochastic backaction term, $\mathcal{H}[\hat{M}]\rho=\hat{M}\rho+\rho\hat{M}^{\dagger}-\operatorname{Tr}\left[\left(\hat{M}+\hat{M}^{\dagger}\right)\rho\right]\rho$, describes continuous monitoring of an observable $\hat{M}$, with $\eta$ the measurement efficiency ($0\leq\eta\leq 1$). The Wiener increment, $dW_{t}$, captures the randomness inherent in quantum measurement outcomes.
@@ B.2.1 Setup for Parameteric Oscillator
$\hat{L}_{1}=\sqrt{\gamma\left(\bar{n}_{{th}}+1\right)}\hat{{a}},\quad\hat{L}_{2}=\sqrt{\gamma\bar{n}_{\mathrm{th}}}\hat{a}^{\dagger},$ (8)
@@
We employ heterodyne detection of the field operator $\hat{a}$. Based on this measurement scheme, we can get seperate the real and imaginary parts to obtain quadrature values that roughly correspond to $X$ and $P$ while adding quantum noise and measurement uncertainities due to quantum mechanical effect [46].
In the following Tables 3(b) & 3(a), we describe the parameters used to generate our dataset by solving the SME in order to generate training data.
Parameter Value
@@
thermal occup. $\langle n_{th}\rangle$ $\mathrm{Uniform}(0.1,0.5)$
@@
(a) In-distribution parameters. Time step $dt=0.5$, total duration $T=600$.
Parameter Value
@@
thermal occup. $\langle n_{th}\rangle$ $\mathrm{Uniform}(0.6,0.7)$
@@
(b) Out-of-distribution parameters. Time step $dt=0.5$, total duration $T=600$.
Table 3: Parameter ranges used in training and evaluation of quantum trajectories. The in-distribution set represents conditions seen during training, while the out-of-distribution set introduces significant shifts in frequency, squeezing, thermal noise, and measurement efficiency.